OpenStudy (gabbyalicorn):

The diameter of the larger circle is 20.4 cm. The diameter of the smaller circle is 10.8 cm. What is the approximate area of the shaded region? A. 326.7 cm² B. 289.4 cm² C. 235.1 cm² D. 72.3 cm²

OpenStudy (anonymous):

\[Area~of~larger~circle-Area~of~smaller~circle\]

OpenStudy (anonymous):

find the area of both circles, then subtract the small circle form the big one

OpenStudy (gabbyalicorn):

oh i kinda maybe sorta get this 1 so 20.4 x 3.14 10.8 x 3.14 then u sutract the 2

OpenStudy (anonymous):

Do you remember the area of a circle formula? We've gone over it twice now

OpenStudy (gabbyalicorn):

64.056 - 33.912 and yes A= π x r ^2

OpenStudy (anonymous):

And you know that \[\Large Radius=\large \frac{1}{2}Diameter\] right?

OpenStudy (anonymous):

\[\sqrt{?}\]

OpenStudy (gabbyalicorn):

right!

OpenStudy (gabbyalicorn):

√ = radius= 1/2 diameter

OpenStudy (anonymous):

\[\sqrt{}=\color{aqua}r\color{red}a\color{blue}d\color{green}i\color{yellow}u\color{brown}s?\]

OpenStudy (gabbyalicorn):

long story short i was saying right but witch a check mark ^u^

OpenStudy (anonymous):

Oh. I was wondering what sqrt's had to do with it.

OpenStudy (gabbyalicorn):

sorry, for the confusion

OpenStudy (anonymous):

Anyways, the questions give us the diameters. We have to halve those to get the radius so that we can plug into the area of a circle equation.

OpenStudy (anonymous):

\[(\pi*\frac{20.4}{2}^{2})-(\pi*\frac{10.8}{2}^{2})\] is our overall equation for this problem.

OpenStudy (anonymous):

Divide before squaring though

OpenStudy (gabbyalicorn):

653.3712 - 183.1248 = 470.2469

OpenStudy (gabbyalicorn):

?

OpenStudy (anonymous):

Divide BEFORE squaring. 20.4/2 = 10.2. 10.2^2=?

OpenStudy (gabbyalicorn):

aww man

OpenStudy (gabbyalicorn):

104.04

OpenStudy (anonymous):

104.04*pi=?

OpenStudy (gabbyalicorn):

326.6856

OpenStudy (anonymous):

Now repeat the process for 10.8

OpenStudy (gabbyalicorn):

366.2496

OpenStudy (anonymous):

10.8/2=?

OpenStudy (gabbyalicorn):

oh 5.4

OpenStudy (anonymous):

5.4^2=?

OpenStudy (gabbyalicorn):

29.16

OpenStudy (anonymous):

29.16*pi=?

OpenStudy (gabbyalicorn):

91.5624

OpenStudy (anonymous):

now subtract

OpenStudy (gabbyalicorn):

so 91.5624 - 29.16 = 62.4024

OpenStudy (gabbyalicorn):

waot no!!!

OpenStudy (anonymous):

Come on man. Look at my very first reply in the thread

OpenStudy (gabbyalicorn):

I was wrong sorry

OpenStudy (gabbyalicorn):

91.5624 -366.2496

OpenStudy (anonymous):

326-91 bruh

OpenStudy (gabbyalicorn):

oh :/ sorry. 235 C

OpenStudy (anonymous):

yeah that'd be it

OpenStudy (gabbyalicorn):

Thank you!

OpenStudy (anonymous):

No problem :)